OpenStudy (xxlyh):

Help please!! Will Fan and Medal Two lines, A and B, are represented by the equations given below: Line A: x + y = 6 Line B: x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There is no solution. It is (6, 4). It is (4, 6).

OpenStudy (anthonyym):

You can equate 6 = 4. Since 6 = x+y and 4 = x+y, then 6 = 4. 6 does not equal 4, so there is no solution.

OpenStudy (xxlyh):

Thank you!! I was leaning towards that but I just wanted to make sure xD

OpenStudy (anthonyym):

You could have also converted the lines to slope intercept form and saw that they have the same slope and there will be no intersection (solution).

OpenStudy (anthonyym):

No problem!

OpenStudy (xxlyh):

do you mind helping me with one more? I'm a little confused with it

OpenStudy (anthonyym):

Yes sure, if it's a different question just post it as a new question.